OpenStudy (anonymous):

write an expression for dy/dx if x^2 - 3xy + 2y^2 = 10

OpenStudy (lgbasallote):

@sami-21 ...know what he means?

OpenStudy (anonymous):

he meant to differentiate it implicitly .

OpenStudy (lgbasallote):

hmm...then isolate dy/dx?

OpenStudy (anonymous):

yes.

OpenStudy (lgbasallote):

sounds doable

OpenStudy (anonymous):

ok differentiate both sides with respect to x use chain and product rules. (derivative of constant is zero. 2x-3(y+xdy/dx)+4ydy/dx=0 solve this for dy/dx. can you do this ??

OpenStudy (anonymous):

im trying

OpenStudy (anonymous):

not getting anywhere

OpenStudy (anonymous):

2x-3y-3xdy/dx+4ydy/dx=0 (4y-3x)dy/dx=3y-2x is it ok now ???

OpenStudy (anonymous):

yes.. thanks alot